Find the equation of a line using the point (7,2) and slope (-4/5)
Answer by TimothyLamb(4379)

slope = m
m = dy/dx
dy = m*dx
---
y - 2 = (-4/5)(x - 7)
y = (-4/5)(x - 7) + 2
y = (-4/5)x + 28/5 + 2
y = (-4/5)x + 28/5 + 10/5
---
answer:
y = (-4/5)x + 38/5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php